OpenStudy (anonymous):

which of the following id the pyramidal number sequence? A. 1, 2, 4, 8, 16, ... B. 1, 4, 9, 16, 25, ... C. 1, 5, 14, 30, 55, ... D. 1, 1, 2, 3, 5, ... E. 1, 3, 6, 10, 15, ..

OpenStudy (anonymous):

* Which of the following is the pyramidal number sequence?

OpenStudy (antiprime):

it isnt d

OpenStudy (antiprime):

its b

OpenStudy (antiprime):

Let S(n) be the nth square pyramidal number. S(n) is defined by the expression S(n) = 1 + 4 + 9 + ... + (n-1)² + n². n terms To find an explicit formula for this sum, consider the sum of n consecutive cubes from 2³ to (n+1)³, that is, the sum Σj=1...n(1+j)³. Σj=1...n(1+j)³ = Σj=1...n(1 + 3j + 3j² + j³) Σ(1+j)³ - Σj³ = Σ1 + 3Σj + 3Σj² (1+j)³ - 1 = Σ1 + 3Σj + 3Σj² (n+1)³ - 1 = n + 3n(n+1)/2 + 3Σj² (n+1)³ - 1 - n - 3n(n+1)/2 = 3Σj² n³ + 1.5n² + 0.5n = 3Σj² n(n+1)(2n+1)/6 = Σj² Thus, S(n) = n(n+1)(2n+1)/6.

OpenStudy (antiprime):

does this help @sunnyday88

OpenStudy (antiprime):

@sunnyday88 @sunnyday88 @sunnyday88

OpenStudy (antiprime):

@KendrickLamar2014

OpenStudy (antiprime):

@pooja195

OpenStudy (antiprime):

@jigglypuff314

OpenStudy (antiprime):

its b

OpenStudy (antiprime):

its b Let S(n) be the nth square pyramidal number. S(n) is defined by the expression S(n) = 1 + 4 + 9 + ... + (n-1)² + n². n terms To find an explicit formula for this sum, consider the sum of n consecutive cubes from 2³ to (n+1)³, that is, the sum Σj=1...n(1+j)³.

OpenStudy (antiprime):

@mathmale

OpenStudy (anonymous):

It was C but Thanks :)